OpenStudy (anonymous):

A particle is in equalibrium under the action of three forces 5 N at 53 above +ve x axis and 10N at 37 below -ve x axis find the magnitude of 3rd force?

OpenStudy (anonymous):

i think..it is since it is not given in the question

OpenStudy (anonymous):

5 cos37 i + 5sin 37 j = 4i + 3 j

OpenStudy (anonymous):

10 cos 37 i + 10 sin 37 j = 8 i + 6j

OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

Can u correct me..

OpenStudy (anonymous):

Find the resultant?

OpenStudy (anonymous):

10 cos 37 (-i) + 10 sin 37 (-j) = 8 i - 6j

OpenStudy (anonymous):

\[\sqrt{a^2 + b^2 + 2abcosx}\]

OpenStudy (anonymous):

i got..it.....

OpenStudy (anonymous):

\[\sqrt{125} = 5\sqrt{5}\]

OpenStudy (anonymous):

is this correct..)

OpenStudy (anonymous):